OpenStudy (anonymous):

Which of th following is the inverse of f(x)= 2-x over 5 ?

OpenStudy (cwrw238):

f(x) = 2-x / 5 2 - x = 5f(x) x = 2 - 5 f(x) inverse f-1 x = 2 - 5x

OpenStudy (anonymous):

good Question. :)